y"+(y') ^2=2e^-y find the general solution
1
Expert's answer
2020-10-26T13:31:34-0400

"\displaystyle

y'' + (y')^2 = 2e^{-y}\\

\textsf{Let}\, y = ln(f(x))\\

y' = \frac{f'(x)}{f(x)}\\

y'' = \frac{f(x)f"(x) - f'(x)\u00b2}{f(x)\u00b2}\\\\\n\ny'' + (y')^2 = 2e^{-y} \\implies \\frac{f(x)f"(x) - f'(x)^2}{f(x)²} + \left(\frac{f'(x)}{f(x)}\right)^2 = \frac{2}{f(x)} \\

\frac{f(x)f"(x)}{f(x)\u00b2} - \\frac{2}{f(x)} = 0\\\\\n\n\n\\frac{f"(x)}{f(x)} - \frac{2}{f(x)} = 0\\


\implies f"(x) - 2 = 0\\\\\n\n\nf"(x) = 2\\

\implies \int f"(x)\\, \\mathrm{d}x = \\int 2 \\, \\mathrm{d}x\\\\\n\nf'(x) = 2x + C_1\\\\\n\n\n\\int f'(x)\\, \\mathrm{d}x = \\int 2x + C_1\\, \\mathrm{d}x\\\\\n\nf(x) = x^2 + C_1x + C_2\\\\\n\n\n\\therefore y = \\ln(x^2 + C_1x + C_2)\\\\\\textsf{is the general solution}\\\\\n\\textsf{to the differential equation, where}\\\\ C_1 \\, \\textsf{and}\\, C_2 \\textsf{are arbitrary constants.}"